OpenStudy (anonymous):

Could somebody help me identify 3 graphs? I'll a medal out!

OpenStudy (anonymous):

y=1/x-4 -2

OpenStudy (anonymous):

y=1/x+2 +4

OpenStudy (anonymous):

@jk67 @Isaiah.Feynman

OpenStudy (anonymous):

y=1/x-2 -3

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

These all follow the same basic layout, y = 1/(x-k) + m. You know there will be a vertical asymptote at k, and the whole thing will be moved up by m. For the first question, the vertical asymptote will be at +4 because its 1/(x-4), and the horizontal asymptote will be at -2 because of the -2 at the end. 3c and 3d are the same picture to me, so it's one of those.